# FullLoRA-AT: Efficiently Boosting the Robustness of Pretrained Vision Transformers

### v1

@@ 3.2.1 LNLoRA
To address these limitations, we introduce the LNLoRA module, which incorporates an additional layernorm preceding the existing LoRA module. The purpose of LNLoRA is to learn the differences in feature magnitude between standard and adversarial trained models via the extra layernorm, allowing the LNLoRA module to better approximate the differences between the parameters of the two models. Consider an image feature represented as $x$, undergoing an initial matrix transformation $h=Wx$. With the addition of our proposed layernorm-included LNLoRA module, it yields:
$h=Wx+BA\cdot\textrm{LN}(x)=Wx+BA(\alpha x+\beta),$ (5)
@@
where $\alpha$ and $\beta$ denote the scale and bias coefficients in the layernorm. The integration of layernorm into the LNLoRA module significantly enhances the model’s ability to adapt to adversarial examples, thereby boosting its robustness with a small number of learnable parameters efficiently.
@@
Table 1: Ablation study on the use of LoRA/LNLoRA in different components of the ViT model.
@@
Patch Embedding MSA block MLP block Classification Head LoRA/LNLoRA Paras (M) (%) (%) (%) (%) (%) (h)
@@
LoRA [20] $\surd$ $\surd$ LoRA 4.72 84.11 47.14 49.96 49.74 45.13 8.10
(a) $\surd$ $\surd$ $\surd$ LoRA 4.74 84.36 47.53 50.31 50.20 45.59 8.17
(b) $\surd$ $\surd$ $\surd$ LoRA 4.74 84.21 47.68 50.49 50.31 45.67 8.18
(c) $\surd$ $\surd$ $\surd$ $\surd$ LoRA 4.77 85.55 48.75 50.83 50.43 46.26 8.23
FullLoRA-AT (ours) $\surd$ $\surd$ $\surd$ $\surd$ LNLoRA 4.85 87.12 49.76 51.38 50.83 46.94 8.38
3.2.2 FullLoRA-AT
Based on the aforementioned LNLoRA module, we introduce the FullLoRA-AT method, a novel and lightweight approach designed to enhance the adversarial robustness of standardly pretrained models in a parameter-efficient manner. As depicted in Fig. 1, the LNLoRA module is seamlessly integrated into several key components of the ViT model, including the patch embedding layer, MSA block, MLP block, and the classification head.
Traditional LoRA methods predominantly focus on modifying parameters within the MSA and MLP blocks of the ViT model, while neglecting the parameters of other model components. However, given the significant differences in image features between clean and adversarial examples, it becomes essential to adaptively finetune the low-level layers of the model, which may facilitate the model’s ability to effectively learn features from adversarial examples that are crucial for accurate classification. Therefore, we extend the LNLoRA module to also encompass the patch embedding layer of the ViT model.
Additionally, given the critical role of parameters in the classification head layer for determining classification outcomes, and the usual practice of unfreezing this layer during finetuning, we also apply the LNLoRA module to the classifier head layer. This extension aims to facilitate a more parameter-efficient adaptation of the model when confronting adversarial examples. Our subsequent ablation studies in the experiment further reveal that the integration of LNLoRA across these key components of ViT model, as opposed to the traditional LoRA method, clearly further boosts the robustness of the model against adversarial attacks.
It is worth mentioning that our FullLoRA-AT method operates independently of existing advanced adversarial training techniques. This characteristic allows for seamless integration with others, leading to further improvement in the efficiency and performance of adversarial finetuning.
4 Experiment
In this section, we first introduce the experimental setting, including datasets, models, comparison methods, and evaluation metrics. Subsequently, we conduct the ablation study to explore the impact of employing LoRA and LNLoRA modules on different components of the ViT model on model robustness. This is followed by comparisons with existing methods on various models and datasets. We then analyze the impact of using different rank $r$ values on the effectiveness of FullLoRA adversarial training. Finally, we integrate FullLoRA-AT with other advanced methods to further accelerate the adversarial training and improve the robustness of the model, effectively demonstrating the generalization of our approach.
4.1 Experimental Setting
Dataset. We conduct experiments on three commonly used datasets in the field of adversarial training, i.e., CIFAR10 [24], CIFAR100 [24] and Imagenette [19]. CIFAR10 and CIFAR100 each contains images of 10 and 100 classes respectively, with training and test datasets comprising 50,000 and 10,000 images respectively. The image size in these datasets is 32$\times$32. Imagenette, a subset of ImageNet [40], includes 10 classes of images, with training and test datasets consisting of 13,000 and 500 images respectively. The image size in Imagenette is 224$\times$244.
Models. We choose ViT-S [13], ViT-B [13] and Swin-B [31] models to conduct the experiments.
Comparison methods. In our study, we compare our FullLoRA-AT method with other approaches such as LoRA [20], UniPELT [35], and Aurora [52], which are all methods for parameter-efficient finetuning on pretrained models. Additionally, following [38], we also compare our approach against strategies that exclusively update specific components during finetuning, e.g., the patch embedding layer and the scale and offset parameters of all layernorms.
Evaluation metrics. In our evaluation of the robustness of various models against adversarial examples, we employ accuracy metrics under different adversarial attack methods. Specifically, we utilize CW-20 [3], PGD-20, PGD-100 [33], and AutoAttack [9] to evaluate model robustness. CW-k and PGD-k denote CW and PGD attacks with k iterative steps, respectively. The perturbation magnitude for these attacks is set to 8/255, with a step size of 1/255. Additionally, we evaluate the efficiency of finetuning pretrained models for robustness by considering the number of learnable parameters required during finetuning and the time taken for the finetuning process.
Details. In our experiments, finetuning on all datasets is conducted over 40 epochs, with the learning rate being scaled down by a factor of 0.1 after the 35-th and 38-th epochs. In default, we employ standard adversarial training with a PGD-10 attack as our baseline method. The SGD optimizer [39] is utilized for the training process. The learning rate settings vary depending on the finetuning method employed: for fully finetuned models and Aurora [52], a learning rate of 0.1 is used; for LoRA [20] and our FullLoRA-AT, the learning rate is set to 5; for UniPELT [35], the learning rate is 0.5; and for the remaining methods, a learning rate of 1 is applied. In the case of LoRA and FullLoRA-AT, considering the trade-off between model performance and training efficiency, we set the rank $r$ to 32 by default. The experiments are carried out on a single NVIDIA GeForce RTX 3090. For CIFAR10 and CIFAR100 datasets, the batch size is set to 512, and for Imagenette, the batch size is 64.
4.2 Ablation Study
@@
We conduct the ablation study using LoRA and LNLoRA modules on different components of the ViT model, and the results are presented in Tab. 1. The experiment is conducted using the ViT-B model on the CIFAR10 dataset. As clearly shown, compared to the original LoRA method which only adds low-rank matrices in the MSA block and MLP block for finetuning, the application of the LoRA module additionally in the classification head and patch embedding layer respectively (i.e., (a) and (b) in Tab. 1) improves the robustness of model under various attacks. Employing the LoRA module across all four mentioned components (i.e., (c) in Tab. 1) further enhances the robustness of the model. Moreover, replacing the LoRA module with our proposed LNLoRA module can further improve the robustness, demonstrating the greater suitability of our newly proposed LNLoRA module for adversarial finetuning tasks on pretrained models. Additionally, as different modules are incrementally added, not only does the robustness of the model improve, but the clean accuracy of the model for classifying clean images also increases consistently.
4.3 Comparison with Other Methods
We first compare the robustness against different attack methods of pretrained models finetuned using our proposed FullLoRA-AT method with other lightweight finetuning methods on various datasets. The results are shown in Tab. 2. When only the patch embedding layer or the scale and offset parameters of layer normalization are used for adversarial finetuning, the number of trainable parameters is minimal and the training speed is relatively fast, whereas the robustness of the model in this scenario is weak. While Aurora [52], LoRA [20], and UniPELT [35] methods show improvements in model robustness, the Aurora method involves a large number of additional matrix multiplication operations due to CP decomposition of the weight matrices, leading to a longer training time, even exceeding that of full finetuning. Also, the UniPELT method adds a considerable amount of extra parameters, resulting in unsatisfactory training duration.
Our proposed FullLoRA-AT method, at rank $r=32$, our FullLoRA-AT method shows a robustness improvement of 2% $\sim$ 4% on average across different datasets compared to the LoRA method, with similar parameter count and training duration, which clearly demonstrates the effectiveness of our approach. When the rank is decreased to $r=16$, our FullLoRA-AT can still achieves similar or even better model robustness compared to the LoRA method while requiring only half the number of parameters for finetuning and reducing training time. Moreover, compared to fully finetuning the entire model, our method uses only about 5% of the parameters and 75% of the training duration to achieve comparable model robustness. On the dataset of CIFAR-10, our method incurs an average loss of only about 1% in robust accuracy, yet it simultaneously achieves a 4.29% improvement in clean accuracy.
We also conduct experiments on different ViT-based models using the CIFAR10 dataset, and the results are displayed in Tab. 3. Similarly, our method achieves the best results across different models, demonstrating the superiority and generalization of our proposed FullLoRA-AT method.
Table 2: Comparison with other methods across different datasets.
Model Method Trainable Clean Acc CW-20 PGD-20 PGD-100 AutoAttack Time
@@ 4.3 Comparison with Other Methods
CIFAR-10 Fully Finetune 85.15 82.83 50.26 52.85 52.56 48.26 11.17
@@
FullLoRA-AT (ours, r=16) 2.46 85.93 47.68 50.33 49.46 45.82 8.01
FullLoRA-AT (ours, r=32) 4.85 87.12 49.76 51.38 50.83 46.94 8.38
CIFAR-100 Fully Finetune 85.22 61.02 28.61 30.58 30.43 27.49 11.18
Patch Embedding Layer 0.04 26.31 10.67 11.98 11.81 9.51 7.42
Layer Normalization 0.04 34.37 13.16 13.79 13.72 10.84 7.80

### v2

@@ III-B1 LNLoRA
To address these limitations, we introduce the LNLoRA module, which incorporates an additional layernorm preceding the existing LoRA module. The purpose of LNLoRA is to learn the differences in feature magnitude between standard and adversarial trained models via the extra layernorm, allowing the LNLoRA module to better approximate the differences between the parameters of the two models, which can be well validated in Section IV-G. As shown in Fig. 1a, consider an image feature represented as $x$, undergoing an initial matrix transformation $h=Wx$. With the addition of our proposed layernorm-included LNLoRA module, it yields:
$h=Wx+BA\cdot\textrm{LN}(x)=Wx+BA(\alpha\cdot\frac{x-\mu(x)}{\sigma(x)}+\beta),$ (5)
@@
where $\alpha$ and $\beta$ denote the scale and bias coefficients in the layernorm, $\mu(\cdot)$ and $\sigma(\cdot)$ represents the mean and standard deviation of the input. The integration of layernorm into the LNLoRA module significantly enhances the model’s ability to adapt to adversarial examples, thereby boosting its robustness with a small number of learnable parameters efficiently. It should be noted that the introduction of layernorm prevents LNLORA from inheriting the LoRA method’s characteristic of merging parameters into matrix $W$ during inference. However, as demonstrated in subsequent experiments, the additional computational overhead introduced by this design choice is negligible.
III-B2 FullLoRA
Based on the aforementioned LNLoRA module, we introduce the FullLoRA method, a novel and lightweight approach designed to enhance the adversarial robustness of standardly pretrained models in a parameter-efficient manner. As depicted in Fig. 1b, the LNLoRA module is seamlessly integrated into several key components of the ViT model, including the patch embedding layer, MSA block, MLP block, and the classification head.
Traditional LoRA methods predominantly focus on modifying parameters within the MSA and MLP blocks of the ViT model, while neglecting the parameters of other model components. However, given the significant differences in image features between clean and adversarial examples, it becomes essential to adaptively finetune the low-level layers of the model, which may facilitate the model’s ability to effectively learn features from adversarial examples that are crucial for accurate classification. Therefore, we also encompass the LNLoRA module to the patch embedding layer of the ViT model.
Additionally, given the critical role of parameters in the classification head layer for determining classification results, and the usual practice of unfreezing this layer during finetuning, we also apply the LNLoRA module to the classification head layer. This extension aims to facilitate a more parameter-efficient adaptation of the model when confronting adversarial examples. Our subsequent ablation studies in the experiment further reveal that the integration of LNLoRA across these key components of ViT model, as opposed to the traditional LoRA method, clearly further boosts the robustness of the model against adversarial attacks.
It is worth mentioning that our FullLoRA method is independent of existing advanced adversarial training techniques. This characteristic allows for seamless integration with others, leading to further improvement in the efficiency and effectiveness of adversarial finetuning.
IV Experiment
In this section, we first introduce the experimental setting, including datasets, models, comparison methods, and evaluation metrics. Subsequently, we conduct the ablation study to explore the impact of employing LoRA and LNLoRA modules on different components of the ViT model on model robustness. This is followed by comparisons with existing methods on various models and datasets. We then analyze the impact of using different rank $r$ values on the effectiveness of FullLoRA. Finally, we integrate FullLoRA with other advanced methods to further accelerate the adversarial training and improve the robustness of the model, effectively demonstrating the generalization of our approach.
IV-A Experimental Setting
IV-A1 Dataset
We conduct experiments on four commonly used datasets in the field of adversarial training, i.e., CIFAR10 [66], CIFAR100 [66], TinyImageNet [67], Imagenette [68] and ImageNet [46]. CIFAR10 and CIFAR100 each contains images of 10 and 100 classes respectively, with training and test datasets comprising 50,000 and 10,000 images respectively. The image size in these datasets is 32$\times$32. TinyImageNet contains 100,000 images of 200 classes (500 for each class) downsized to 64$\times$64 colored images. Each class has 500 training images, 50 validation images and 50 test images. Imagenette, a subset of ImageNet [46], includes 10 classes of images, with training and test datasets consisting of 13,000 and 500 images respectively. The image size in TinyImageNet, Imagenette and ImageNet is 224$\times$244.
IV-A2 Models
We choose ViT-S [10], ViT-B [10] and Swin-B [12] models to conduct the experiments. In all experiments, we initialize the model with weights pretrained on ImageNet-1K while employing random initialization for the final classification layer.
IV-A3 Comparison Methods
In our study, we compare our FullLoRA method with other approaches such as LoRA [35], UniPELT [36], and Aurora [39], which are all methods for parameter-efficient finetuning on pretrained models. Additionally, following [62], we also compare our approach against strategies that exclusively update specific components during finetuning, e.g., the patch embedding layer and the scale and offset parameters of all layernorms. It should be noted that, in our comparisons, we have combined these methods with adversarial training to enhance the model’s robustness, ensuring a fair comparison.
IV-A4 Evaluation Metrics
In our evaluation of the robustness of various models against adversarial examples, we employ accuracy metrics under different adversarial attack methods. Specifically, we utilize CW-20 [41], PGD-20, PGD-100 [22], and AutoAttack [42] to evaluate model robustness. CW-$k$ and PGD-$k$ denote CW and PGD attacks with $k$ iterative steps, respectively. The perturbation magnitude for these attacks is set to 8/255, with a step size of 1/255. Additionally, we evaluate the efficiency of finetuning pretrained models for robustness by considering the number of learnable parameters required during finetuning and the time taken for the finetuning process.
IV-A5 Details
In our experiments, finetuning on all datasets is conducted over 40 epochs, with the learning rate being scaled down by a factor of 0.1 after the 35-th and 38-th epochs. In default, we employ standard adversarial training with a PGD-10 attack as our baseline method. In our adversarial training, we exclusively use adversarial examples and does not utilize clean images. The SGD optimizer [69] is utilized for the training process. The learning rate settings vary depending on the finetuning method employed: for fully finetuned models and Aurora [39], a learning rate of 0.1 is used; for LoRA [35] and our FullLoRA, the learning rate is set to 5; for UniPELT [36], the learning rate is 0.5; and for the remaining methods, a learning rate of 1 is applied. In the case of LoRA and FullLoRA, considering the trade-off between model performance and training efficiency, we set the rank $r$ to 32 by default. The experiments are carried out on a single NVIDIA GeForce RTX 3090 except ImageNet, which is conducted on four NVIDIA GeForce RTX 3090. In order to ensure the stability and persuasiveness of the experiment, we re-conduct most experiments in our proposed FullLoRA method five times, and show the mean and standard deviation of experimental results.
@@ IV-A5 Details
TABLE I: Ablation study on the use of LoRA/LNLoRA in different components of the ViT model.
@@
Patch Embedding MSA block MLP block Classification Head LNLoRA Paras (M) (%) (%) (%) (%) (%) (h)
@@
LoRA [35] $\surd$ $\surd$ w/o 4.72 84.11 47.14 49.96 49.74 45.13 8.10
(a) $\surd$ $\surd$ $\surd$ w/o 4.74 84.36 47.53 50.31 50.20 45.59 8.17
(b) $\surd$ $\surd$ $\surd$ w/o 4.74 84.21 47.68 50.49 50.31 45.67 8.18
(c) $\surd$ $\surd$ $\surd$ $\surd$ w/o 4.77 85.55 48.75 50.83 50.43 46.26 8.23
FullLoRA (ours) $\surd$ $\surd$ $\surd$ $\surd$ w/ 4.85 87.12 49.76 51.38 50.83 46.94 8.38
IV-B Ablation Study
@@
We conduct the ablation study using LoRA and LNLoRA modules on different components of the ViT model, and the results are presented in Table I. The experiment is conducted using the ViT-B model on the CIFAR10 dataset. As clearly shown, compared to the original LoRA method which only adds low-rank matrices in the MSA block and MLP block for finetuning, the application of the LoRA module additionally in the classification head and patch embedding layer respectively (i.e., (a) and (b) in Table I) improves the robustness of model under various attacks. Employing the LoRA module across all four mentioned components (i.e., (c) in Table I) further enhances the robustness of the model. Moreover, replacing the LoRA module with our proposed LNLoRA module can further improve the robustness, demonstrating the greater suitability of our newly proposed LNLoRA module for adversarial finetuning on pretrained models. Additionally, as different modules are incrementally added, not only does the robustness of the model improve, but also the clean accuracy of the model increases consistently, which further demonstrates the effectiveness of our proposed FullLoRA.
TABLE II: Comparison with other methods across different datasets. The results are obtained using the ViT-B model. The results reported for the FullLoRA method are the mean and standard deviation of repeating the experiment five times.
Dataset Method Trainable Clean Acc CW-20 PGD-20 PGD-100 AutoAttack Time
@@
CIFAR-10 Standard Training 98.65 0.00 0.00 0.00 0.00
Fully Finetune 85.15 82.83 50.26 52.85 52.56 48.26 11.17
@@
FullLoRA (ours) 4.85 87.16$\pm$0.20 50.36$\pm$0.12 51.50$\pm$0.09 51.04$\pm$0.08 46.97$\pm$0.09 8.39$\pm$0.02
Imagenette Standard Training 98.00 0.00 0.20 0.00 0.00
Fully Finetune 85.15 93.40 69.00 68.80 68.60 67.60 8.52
Patch Embedding Layer 0.04 37.60 23.40 25.60 25.60 24.80 5.76
Layer Normalization 0.04 61.40 23.60 28.80 28.80 27.60 5.89
Aurora [39] 0.09 87.20 51.40 53.20 53.00 51.80 9.07
LoRA [35] 4.72 94.00 63.80 62.00 61.20 60.00 6.15
UniPELT [36] 11.42 91.40 58.00 57.80 56.20 55.40 9.58
FullLoRA (ours) 4.85 93.48$\pm$0.23 66.26$\pm$0.26 66.28$\pm$0.23 65.64$\pm$0.17 64.02$\pm$0.11 6.31$\pm$0.02
IV-C Comparison with Other Methods
We first compare the robustness against different attack methods of pretrained models finetuned using our proposed FullLoRA method with other lightweight finetuning methods on various datasets. It should be noted that, in our comparisons, we have combined these methods with adversarial training to enhance the model’s robustness, ensuring a fair comparison. The results are shown in Table II. When only the patch embedding layer or the scale and offset parameters of layer normalization are used for adversarial finetuning, the number of trainable parameters is minimal and the training speed is relatively fast, whereas the robustness of the model in this scenario is weak. While Aurora [39], LoRA [35], and UniPELT [36] methods show improvements in model robustness, the Aurora method involves a large number of additional matrix multiplication operations due to CP decomposition of the weight matrices, leading to a longer training time, even exceeding that of full finetuning. Also, the UniPELT method adds a considerable amount of extra parameters, resulting in unsatisfactory training duration.
Our proposed FullLoRA method shows a robustness improvement of 2% $\sim$ 4% on average across different datasets compared to the LoRA method, with similar parameter count and training duration, which clearly demonstrates the effectiveness of our approach. Moreover, compared to fully finetuning the entire model, our method uses only about 5% of the parameters and 75% of the training duration to achieve comparable model robustness. On the dataset of CIFAR-10, our method incurs an average loss of only about 1% in robust accuracy, yet it simultaneously achieves a 4.29% improvement in clean accuracy.
We also conduct experiments on different ViT-based models using the CIFAR10 dataset, and the results are displayed in Table III. Similarly, our method achieves the best results across different models, demonstrating the superiority and generalization of our proposed FullLoRA method. To further demonstrate the effectiveness of our FullLoRA method, we compare the robustness of different models under the gradient-free attacks (i.e., ZOO [70]) and obfuscated gradient attacks (i.e., BPDA [71]). From Table III, we can clearly see that our proposed FullLoRA method also demonstrates better robustness under the ZOO and BPDA attacks compared to existing methods. This strongly demonstrates the generalization of FullLoRA, showing that it improves the model’s robustness under various attack methods.
TABLE III: Comparison with other methods on different ViT-based models. The results are obtained using the CIFAR10 dataset. The results reported for the FullLoRA method are the mean and standard deviation of repeating the experiment five times.
Model Method Trainable Clean Acc CW-20 PGD-20 PGD-100 AutoAttack BPDA ZOO Time
Paras (M) (%) (%) (%) (%) (%) (%) (%) (h)
ViT-B Standard Training 98.65 0.00 0.00 0.00 0.00 0.03 1.32
Fully Finetune 85.15 82.83 50.26 52.85 52.56 48.26 49.63 61.38 11.17
Aurora [39] 0.09 76.09 38.50 41.85 41.67 39.87 37.21 49.52 11.48
LoRA [35] 4.72 84.11 47.14 49.96 49.74 45.13 45.71 56.79 8.10
UniPELT [36] 11.42 84.63 47.63 49.49 48.99 44.63 45.32 55.17 13.11
FullLoRA (ours) 4.85 87.16$\pm$0.20 50.36$\pm$0.12 51.50$\pm$0.09 51.04$\pm$0.08 46.97$\pm$0.09 47.65$\pm$0.27 59.46$\pm$0.33 8.39$\pm$0.02
ViT-S Standard Training 96.09 0.00 0.00 0.00 0.00 0.03 1.74
Fully Finetune 47.34 79.94 48.02 50.84 50.62 46.22 46.92 58.34 5.31
Aurora [39] 0.08 70.52 36.12 38.91 38.79 35.24 36.24 44.87 5.46
LoRA [35] 3.15 78.31 44.87 46.10 45.98 41.37 42.67 52.68 3.87
UniPELT [36] 7.75 73.40 39.82 43.68 43.49 40.09 41.38 50.96 5.93
FullLoRA (ours) 3.25 78.97$\pm$0.17 45.69$\pm$0.10 48.57$\pm$0.10 48.41$\pm$0.07 44.95$\pm$0.06 45.19$\pm$0.31 55.83$\pm$0.29 3.92$\pm$0.01
Swin-B Standard Training 97.88 0.00 0.00 0.00 0.00 0.02 1.16
Fully Finetune 86.69 81.29 48.36 50.99 50.88 46.57 46.98 58.76 6.44
Aurora [39] 0.09 75.13 37.52 38.17 38.03 35.26 36.61 45.13 4.79
LoRA [35] 6.16 80.94 45.32 46.50 46.37 43.08 43.84 54.62 4.66
UniPELT [36] 11.58 79.55 44.16 45.28 44.93 41.39 42.42 53.55 5.07
FullLoRA (ours) 6.29 83.66$\pm$0.19 46.78$\pm$0.11 48.23$\pm$0.11 48.13$\pm$0.12 44.21$\pm$0.09 45.91$\pm$0.21 56.97$\pm$0.17 4.81$\pm$0.01
IV-D Experiments on More Datasets
We further conduct experiments on the CIFAR-100, TinyImageNet, and ImageNet datasets. The results in Table IV clearly show that our method also achieves the best results on these three datasets, fully demonstrating the good generalizability of the approach we propose. On the CIFAR-100 and TinyImageNet datasets, our proposed FullLoRA method has shown an average robustness improvement of over 3% compared to LoRA under various attacks, effectively showcasing the superiority of our method. On the more challenging ImageNet dataset, due to training costs and time constraints, we only compare the results with the LoRA method. From the table, it can be seen that our FullLoRA also achieves an average robustness improvement of over 2% compared to the LoRA method.
TABLE IV: More results on the dataset of CIFAR-100, TinyImageNet and ImageNet. The results reported for the FullLoRA method are the mean and standard deviation of repeating the experiment five times.
Dataset Method Trainable Clean Acc CW-20 PGD-20 PGD-100 AutoAttack Time
Paras (M) (%) (%) (%) (%) (%) (h)
CIFAR-100 Standard Training 78.39 0.00 0.00 0.00 0.00
Fully Finetune 85.22 61.02 28.61 30.58 30.43 27.49 11.18